OpenStudy (anonymous):

A trapezoid has an area given by A = (b 1 + b 2)h, where b 1 is the length of one base, b 2 is the length of the other base, and h is the height. If the trapezoid has an area of 14 square units, solve for b 1 in terms of the other variables.

OpenStudy (anonymous):

Can you help me?

OpenStudy (campbell_st):

so so do the opposite to both sides of the equation... and are you sure the A = (b1 + b2)h is correct?

OpenStudy (anonymous):

That is what the question said.

OpenStudy (anonymous):

It's supposed to have 1/2 in front of it. So A = 1/2 (b1+b2)h

OpenStudy (campbell_st):

ok... thats what I was thinking ok so you know 14 = 1/2(b1 + b2)h so the process is doing the opposite of the right hand side to the left side what is the opposite of dividing by 2?

OpenStudy (anonymous):

Multiplying by 2.

OpenStudy (campbell_st):

great that gets rid of the half so 28 = (b1 + b2)h ok so now you need to opposite of multiplying by h

OpenStudy (anonymous):

So divide by h which would then make it 28 = b1+b2/h?

OpenStudy (campbell_st):

well nearly its 28/h = b1 + b2 what ever you do... it has to happen on both sides of the equation on the right h/h = 1 ok... last step... the opposite of adding b2..?

OpenStudy (anonymous):

subtracting b2. So 28/h - b2 = b1

OpenStudy (campbell_st):

thats correct... well done

OpenStudy (campbell_st):

or b1 = 28/h - b2 both are correct

OpenStudy (anonymous):

Thank you!